# Enumerating log rational curves on some toric varieties

### v2

@@ 5.2. Logarithmic Tevelev degrees of $\mathrm{Bl}_{[0:1:0],[0:0:1]}({\mathbb{P}}^{2})$
[CIL24, Conjecture 15] predicts that $\mathrm{logTev}^{X}_{\Gamma}$ either equals 0 or the value of the integral in Proposition 5.2.1(vi). Statements (i)-(iv) and (vi) of Proposition 5.2.1 show that $\mathrm{logTev}^{X}_{\Gamma}$ does in fact equal the right hand side of (vi) if
@@
This value is non-zero and differs from the value $24\cdot(5\cdot 5)\cdot\binom{3}{1}\cdot\binom{3}{1}$ of the integral in Proposition 5.2.1(vi). In particular, [CIL24, Conjecture 15] does not hold.
@@
Thus, $Z$ consists of $24\cdot 5$ components, each isomorphic to ${\mathbb{P}}^{1}_{4,1}\times{\mathbb{P}}^{1}_{5,1}$; in Lemma 5.2.3 below, we show that this is true as a scheme. More precisely, each copy of ${\mathbb{P}}^{1}_{4,1}\times{\mathbb{P}}^{1}_{5,1}$ is viewed as a subscheme of $Q_{\Gamma}(X)$ via the section cut out by the equations $g_{4}=g_{5}=0$.
@@
The normal bundle $N_{Z_{\bullet}/Q_{\Gamma}(X)}$ has a rank 6 trivial sub-bundle $N_{{\mathbb{P}}^{1}_{4,1}\times{\mathbb{P}}^{1}_{5,1}/\mathcal{B}}$. The quotient is cut out by the vanishing of $g_{4}\in H^{0}({\mathbb{P}}^{1},\mathcal{O}(6)(-D_{1}-D_{4}))$ and $g_{5}\in H^{0}({\mathbb{P}}^{1},\mathcal{O}(6)(-D_{2}-D_{5}))$. Therefore,
@@
[[[[ $(1+\gamma_{3}\epsilon)(z-p_{5}-\gamma_{11}\epsilon)(z-p_{5}-\gamma_{21}\epsilon)(z-p_{1}-\gamma_{31}\epsilon)(z-p_{2}-\gamma_{32}\epsilon)(z-p_{3}-\gamma_{33}\epsilon)(z-p_{4}-\gamma_{34}\epsilon):$

### v3

@@ 5.2. Logarithmic Tevelev degrees of $\mathrm{Bl}_{[0:1:0],[0:0:1]}({\mathbb{P}}^{2})$
[cil, Conjecture 15] predicts that $\mathrm{logTev}^{X}_{\Gamma}$ either equals 0 or the value of the integral in Proposition 5.2.1(vi). Statements (i)-(iv) and (vi) of Proposition 5.2.1 show that $\mathrm{logTev}^{X}_{\Gamma}$ does in fact equal the right hand side of (vi) if
@@
This value is non-zero and differs from the value $24\cdot(5\cdot 5)\cdot\binom{3}{1}\cdot\binom{3}{1}$ of the integral in Proposition 5.2.1(vi). In particular, [cil, Conjecture 15] does not hold.
@@
Thus, $Z$ consists of $24\cdot 5$ components, each isomorphic to ${\mathbb{P}}^{1}_{4,1}\times{\mathbb{P}}^{1}_{5,1}$; in Lemma 5.2.3 below, we show that this is true as a scheme. More precisely, each copy of ${\mathbb{P}}^{1}_{4,1}\times{\mathbb{P}}^{1}_{5,1}$ is viewed as a subscheme of $Q_{\Gamma}(X)$ via the section $s:\mathcal{B}\hookrightarrow Q_{\Gamma}(X)$ cut out by the equations $g_{4}=g_{5}=0$.
@@
The inclusion $Z_{\bullet}\hookrightarrow Q_{\Gamma}(X)$ factors as
$Z_{\bullet}\hookrightarrow\mathcal{B}\hookrightarrow Q_{\Gamma}(X),$
where $Z_{\bullet}\hookrightarrow\mathcal{B}$ is the inclusion determined by the given ordering of the points $p_{i}$, and $\mathcal{B}\hookrightarrow Q_{\Gamma}(X)$ is the section $s$ cut out by the equations $g_{4}=g_{5}=0$. The rank 6 sub-bundle $N_{Z_{\bullet}/\mathcal{B}}\subset N_{Z_{\bullet}/Q_{\Gamma}(X)}$ is trivial, because it is given by an inclusion $({\mathbb{P}}^{1})^{2}\subset({\mathbb{P}}^{1})^{8}$, which is pulled back from an inclusion of a point in $({\mathbb{P}}^{1})^{6}$. The quotient is cut out by the vanishing of $g_{4}\in H^{0}({\mathbb{P}}^{1},\mathcal{O}(6)(-D_{1}-D_{4}))$ and $g_{5}\in H^{0}({\mathbb{P}}^{1},\mathcal{O}(6)(-D_{2}-D_{5}))$. Therefore,
@@
$[$ $(1+\gamma_{3}\epsilon)(z-p_{5}-\gamma_{11}\epsilon)(z-p_{5}-\gamma_{21}\epsilon)(z-p_{1}-\gamma_{31}\epsilon)(z-p_{2}-\gamma_{32}\epsilon)(z-p_{3}-\gamma_{33}\epsilon)(z-p_{4}-\gamma_{34}\epsilon):$